OpenStudy (anonymous):

Evaluate the expression and interpret the work in terms of the context: f(3) Time: 3 Velocity: 96

OpenStudy (anonymous):

what expression?

OpenStudy (anonymous):

I need to find the inverse of f(3). The original function is below that, where the time is 3 and velocity of 96.

OpenStudy (anonymous):

... so, at time = 3, the velocity is 96? ....so f(3) = 96? inverse would be... f'(96) = 3?

OpenStudy (anonymous):

Is that how it works? It's that simple???

OpenStudy (anonymous):

or is it t=f^-1(v)?

OpenStudy (anonymous):

Yes, that is the correct notation. I was being lazy. inverses is just flipping the x and y. (easy answer)

OpenStudy (anonymous):

if I plug in 3 to get 96...then when i find the inverse function (i.e. solve for 3, I need to plug in 96 to get back to 3.)

OpenStudy (anonymous):

or would it be 3^-1(96)?

OpenStudy (anonymous):

f isn't defined (no equation-unless you're hiding it from me.) all we know is when 3 is the input, the output is 96. the inverse f would use the 96 as the input to get 3 as the output. So... f(3) = 96 \[f ^{-1}(96) = 3\]

OpenStudy (anonymous):

Oh, the equation might b3e v=f(t)=32t so the inverse of that would be t=f^-1(v)

OpenStudy (anonymous):

no. v = 32t (swap v and t) t = 32v (solve for v) t/32 = v therefore, \[f ^{-1}(t) = t/32\]

OpenStudy (anonymous):

ok thanks!